# Exploring Holomorphic Retracts

### v1

@@ Conclusion
As mentioned in the introduction until this stage, we have for the most part, restricted attention to retracts of bounded domains; though the domains in the previous section allows topologically non-triviality, the boundedness assumption prevailed. To go beyond, to more general settings, we require the following lemma:
@@
Next we shall detail the proof of retraction mappings of ×$\mathbb{C}\times\Delta$.
@@
Equating degree $2$ terms on both sides, we get $P_{2}(P_{1},0)=0$. This implies that $P_{2}(z,w)=w(Q_{1}(z,w))$, where $Q_{1}(z,w)$ is a polynomial in $z$ and $w$. Substituting $P_{2}$ in 8.2, we get
@@
Proof of theorem 1.17.
@@
We shall next record the proof of theorem 1.16, whose verifying details will be shortened as most of its arguments are contained in the prequel. To explicitly spell out the special case of domains contained in Euclidean space: taking $X$ to be the simplest non-compact Riemann surface namely $\mathbb{C}$, we have a description of all the retracts of ×$\mathbb{C}\times D$ where $D$ is any hyperbolic domain in $\mathbb{C}^{N}$.
@@
We first give an example of a holomorphic retract of $\mathbb{C}^{2}=\mathbb{C}\times\mathbb{C}$ which can neither be realized as the graph over the first factor nor the second. This is discussed in the following
@@
is a holomorphic retract of $\mathbb{C}^{2}$ that can neither be written as a graph of any holomorphic map on the first factor nor the second; indeed, it is the image of the holomorphic retraction map $f$ on $\mathbb{C}^{2}$ defined by $f(z,w)=(e^{zw},zw~{}e^{-zw})$ and $Z$ can in-fact not be written as a graph of any holomorphic map on any one-dimensional linear subspace of $\mathbb{C}^{2}$.
@@
Let $L_{1}$ be an arbitrary ray emanating from origin. For each $z\in\mathbb{C}$, there exists an open disk $D$ such that $D\subset\mathbb{C}\setminus L_{1}$. Let $\ell_{1}$ denotes the branch of logarithm on $\mathbb{C}\setminus L_{1}$. Choose a ray $L_{2}$ starting from origin such that $L_{2}$ intersects $D$ non-vacuously. $L_{2}$ divides $D$ into two regions namely $D^{+}~{}\text{and}~{}D^{-}$. Let $\ell_{2}$ denotes the branch of logarithm on $\mathbb{C}\setminus L_{2}$.
Since $h(w)^{w}=g(w)$ for all $w\in\mathbb{C}$, $w\ell_{1}(h(w))=\ell_{1}(g(w))~{}\text{for all}~{}w\in D^{+}\cup D^{-}$ and this implies that $h(w)=e^{\frac{\ell_{1}(g(w))}{w}}$. Similarly, we get $h(w)=e^{\frac{\ell_{2}(g(w))}{w}}~{}\text{for all}~{}w\in D^{+}\cup D^{-}$.
On $D^{+}\cup D^{-}$,
$e^{\frac{\ell_{1}(g(w))}{w}}=e^{\frac{\ell_{2}(g(w))}{w}}$
@@
(9.1) $\ell_{1}(g(w))-\ell_{2}(g(w))=2n\pi iw~{}$
But $\ell_{1}(g(w))-\ell_{2}(g(w))$ is an integer multiple of $2\pi i$ for any $w\in D^{+}\cup D^{-}$. So, (9.1) holds only for $n=0$. Hence we get $\ell_{1}(g(w))=\ell_{2}(g(w))~{}\text{for all}~{}w\in D^{+}\cup D^{-}$. This implies that $\ell_{2}$ is a branch of logarithm on $\mathbb{C}^{*}$. This contradicts the fact that there does not exists a branch of logarithm on $\mathbb{C}^{*}$. ∎
We conclude by reporting some results about polynomial retracts of $\mathbb{C}^{2}$, the simplest of domains admitting non-trivial retracts. It seems difficult to get a complete characterization of holomorphic retracts of $\mathbb{C}^{2}$. However it is possible to deduce a complete characterization of polynomial retracts on $\mathbb{C}^{2}$ from known results about the polynomial ring in complex two variables. Even this deduction is not entirely trivial and needs some work which we record here. In order to lay this work out properly, we begin with some definitions.
@@
Let $K[x,y]$ be the polynomial algebra in two variables over a field $K$ of characteristic 00. A subalgebra $R$ of $K[x,y]$ is called a retract if it satisfies any of the following equivalent conditions:
  1. (R1)
    There is an idempotent homomorphism (a retraction, or projection) $\varphi:K[x,y]\rightarrow$ $K[x,y]$ such that $\varphi(K[x,y])=R$.
  2. (R2)
    There is a homomorphism $\varphi:K[x,y]\rightarrow R$ that fixes every element of $R$.
  3. (R3)
    $K[x,y]=R\oplus I$ for some ideal $I$ of the algebra $K[x,y]$.
  4. (R4)
    $K[x,y]$ is a projective extension of $R$ in the category of $K$-algebras. In other words, there is a splitting exact sequence $1\rightarrow I\rightarrow K[x,y]\rightarrow R\rightarrow 1$, where $I$ is the same ideal as in (R3) above.
To provide just one concrete example to illustrate the above definitions, take $p=x+x^{2}y$; then $K[p]$ is a retract of $K[x,y]$ (the retraction map here, is given by $x\mapsto p$ , $y\mapsto 0$).
In [Cost], Costa proved that every proper retract of $K[x,y]$ has the form $K[p]$ for some polynomial $p\in K[x,y]$.
Next we shall show how polynomial retraction maps on $\mathbb{C}^{2}$ transforms to a retraction map on the polynomial ring $\mathbb{C}[z,w]$. We can achieve this using the following lemma.
@@
A holomorphic map $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by $\rho(z,w)=(p(z,w),q(z,w))$, is a polynomial retraction if and only if there exists a retraction map $\tilde{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ whose components are polynomials $p(z,w)$ and $q(z,w)$.
@@
Assume that $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ is a retraction map on $\mathbb{C}^{2}$ such that whose components are polynomials namely $p(z,w)$ and $q(z,w)$. Define $\tilde{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ defined by $z\mapsto p(z,w),~{}w\mapsto q(z,w)$. Since $\rho\circ\rho=\rho$, we get $p(p(z,w),q(z,w))=p(z,w)$ and $q(p(z,w),q(z,w))=q(z,w)$. This imply that $\tilde{\rho}\circ\tilde{\rho}=\tilde{\rho}$. So, we get that $\tilde{\rho}$ is a retraction on $\mathbb{C}[z,w]$ whose image is precisely $R$, the subalgebra generated by $p,q\in\mathbb{C}[z,w]$.
Conversely, assume $\tilde{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ is a retraction map. Then there exists $p(z,w),\\ q(z,w)\in\mathbb{C}[z,w]$ such that $\rho(z)=p(z,w)$ and $\rho(w)=q(z,w)$. Note that the retract $R$ is the subalgebra generated by $p(z,w),~{}q(z,w)$.
Define $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by $\rho(z,w)=(p(z,w),q(z,w))$. Since $p,q\in R$, $\tilde{\rho}(p)=p$ and $\tilde{\rho}(q)=q$. This implies that $p(p(z,w),q(z,w))=p(z,w)$ and $q(p(z,w),q(z,w))=q(z,w)$. Hence we get that $\rho$ is a polynomial retraction on $\mathbb{C}^{2}$.
@@
Let $Z$ be a non-trivial polynomial retract of $\mathbb{C}^{2}$. There exists a retraction map $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ is of the form $\rho(z,w)=(p(z,w),q(z,w))$, where $p,q\in\mathbb{C}[z,w]$ such that $\rho(\mathbb{C}^{2})=Z$. By lemma 9.4, $\rho$ induces a retraction map $\tilde{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ such that $\tilde{\rho(z)}=p(z,w)$ and $\tilde{\rho}(w)=q(z,w))$, where $p,q\in\mathbb{C}[z,w]$. Note that ${\rm image}(\tilde{\rho})=\mathbb{C}[p,q]$, the subalgebra generated by $p,q$.
By Proposition 3.1 in [Cost], $\mathbb{C}[p,q]=\mathbb{C}[r]$, where $r=r(z,w)\in\mathbb{C}[z,w]$. Write $p=q_{1}(r),~{}q=q_{2}(r)$, where $q_{1},q_{2}\in\mathbb{C}[z]$ So $\rho(z,w)=(q_{1}(r(z,w)),q_{2}(r(z,w)))$.
Suppose $q_{1}$ and $q_{2}$ have degree $n\geq 1$ and $m\geq 1$ respectively. By Abhyankar - Moh theorem in [Abynkr], either $n$ divides $m$ or $m$ divides $n$. Suppose $deg(q_{1})=k\cdot deg(q_{2})$ for some integer $k\geq 1$.
Consider an automorphism $\psi:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by $\psi(z,w)=(z-cw^{k},w)$, where $c$ is chosen so that leading term of polynomial $q_{1}-cq_{2}^{k}$ cancel out.
@@
Continuing this process, we eventually arrive at a pair of polynomials. One of which is a constant polynomial $c$, and the other polynomial which is denoted by $h$, then we must have $\mathbb{C}[h]=\mathbb{C}[r]$ (i.e. $h=c_{1}r+c_{2}$). Finally we get $\phi\in Aut(\mathbb{C}^{2})$ such that
$\phi\circ\rho\circ\phi^{-1}(z,w)=(c_{1}r(z,w)+c_{2},c).$
After applying conjugation of affine automorphism, we get $\phi(z,w)$ is of the form $(s(z,w),0)$, where $s(z,w)\in\mathbb{C}[z,w]$ and $s(0,0)=0$. Since $\phi\circ\rho\circ\phi^{-1}$ is a retraction map, we will get $s(s(z,w),0)=s(z,w)$. Write $s(z,0)=a_{1}z+a_{2}z^{2}+\ldots+a_{k}z^{k}$. Note that
$s(s(z,w),0)=a_{1}s(z,w)+\ldots+a_{k}s(z,w)^{k}=s(z,w).$
This implies $a_{1}=1$ and $a_{2}=...=a_{k}=0$. So $s(z,w)=z+w\cdot l(z,w)$, where $l(z,w)\in\mathbb{C}[z,w]$. Hence $Z$ is of the form $\{(\zeta,0):\zeta\in\mathbb{C}\}$.

### v6

@@ Conclusion
As mentioned in the introduction until this stage, we have for the most part, restricted attention to retracts of bounded domains; though the domains in the previous section allows topological non-triviality, as we considered ‘analytic complements’ $D\setminus A$ at the beginning therein, the boundedness assumption prevailed. To go beyond, to more general settings, we require the following lemma:
@@
Next we shall detail the complete determination of the retraction maps of ×$\mathbb{C}\times\Delta$.
@@
Equating degree $2$ terms on both sides, we get $P_{2}(P_{1},0)=0$. This implies that $P_{2}(z,w)=w(Q_{1}(z,w))$, where $Q_{1}(z,w)$ is a polynomial in $z$ and $w$. Substituting $P_{2}$ in 9.2, we get
@@
Next, we turn to the proofs of theorems dealing with domains or manifolds with much greater topological complexity.
Proof of proposition 1.20.
@@
We shall next record the proof of theorem 1.22, whose verifying details will be shortened as most of its arguments are contained in the prequel. To explicitly spell out the special case of domains contained in Euclidean space: taking $X$ to be the simplest non-compact Riemann surface namely $\mathbb{C}$, we have a description of all the retracts of ×$\mathbb{C}\times D$ where $D$ is any hyperbolic domain in $\mathbb{C}^{N}$. Indeed as mentioned earlier, this class contains domains with a wide variety of fundamental (or for that matter higher homotopy and homology) groups.
@@
To begin with a few words in general about the non-trivial (holomorphic) retracts of the complex Euclidean space $\mathbb{C}^{N}$ of any given dimension $N$, note that all its retracts are contractible complex submanifolds which are non-compact as $\mathbb{C}^{N}$ does not admit any non-trivial compact submanifold. Apart from being Stein submanifolds, they are non-hyperbolic; indeed, they are ‘Liouville’ submanifolds as their Kobayashi metric vanishes identically (due to lemma 2.25). In-particular, we see as an immediate consequence of the uniformization theorem that any one-dimensional retract $Z$ of $\mathbb{C}^{N}$ is biholomorphic to $\mathbb{C}$; thereby, $Z\times\mathbb{C}^{N-1}$ is biholomorphic to $\mathbb{C}^{N}$. Thus, we may decompose $\mathbb{C}^{N}$ as $Z\times\mathbb{C}^{N-1}$ as well, though it may be asked as to why we may want such a decomposition. More importantly, our purpose in proving that retracts of the various kinds of bounded balanced domains earlier, are graphs of holomorphic maps, was to cast the retracts in as simple a ‘normal form’ as possible. It is natural to first examine if this as literally true for $\mathbb{C}^{N}$. Towards this, we first give an example of a holomorphic retract of $\mathbb{C}^{2}=\mathbb{C}\times\mathbb{C}$ which can neither be realized as the graph over the first factor nor the second; moreover, it is also not a product of retracts – taking graphs and products of retracts are the two canonical ways in which retracts of product domains arise and this provides an example that cannot be realized as arising directly from either of these ways. This is discussed in the following
@@
is a holomorphic retract of $\mathbb{C}^{2}$ that can neither be written as a graph of any holomorphic map on the first factor nor the second; indeed, it is the image of the holomorphic retraction map $f$ on $\mathbb{C}^{2}$ defined by $f(z,w)=(e^{zw},\;zwe^{-zw})$ and $Z$ can in-fact not be written as a graph of any holomorphic map on any one-dimensional linear subspace of $\mathbb{C}^{2}$.
@@
Let $L_{1}$ be an arbitrary ray emanating from origin and $z$ an arbitrary point of $\mathbb{C}\setminus L_{1}$. Just by the closedness of $L_{1}$ in $\mathbb{C}$, we may pick an open disc $D\subset\mathbb{C}\setminus L_{1}$. As $\mathbb{C}\setminus L_{1}$ is simply connected, there exists atleast one (indeed infinitely many) branch of the (complex) logarithm on $\mathbb{C}\setminus L_{1}$; pick one of them and call it $\ell_{1}$. Choose a ray $L_{2}$ starting from origin such that $L_{2}$ intersects $D$ Note that $L_{2}$ divides $D$ into two regions which shall be referred to by $D^{+}~{}\text{and}~{}D^{-}$. As with $\ell_{1}$, we likewise pick and denote by $\ell_{2}$, a branch of logarithm on $\mathbb{C}\setminus L_{2}$.
Since $h(w)^{w}=g(w)$ for all $w\in\mathbb{C}$, we have $w\ell_{1}(h(w))=\ell_{1}(g(w))~{}\text{for all}~{}w\in D^{+}\cup D^{-}$ which implies that $h(w)=e^{\frac{\ell_{1}(g(w))}{w}}$. Similarly, we get $h(w)=e^{\frac{\ell_{2}(g(w))}{w}}~{}\text{for all}~{}w\in D^{+}\cup D^{-}$.
It then follows that on $D^{+}\cup D^{-}$,
$e^{\frac{\ell_{1}(g(w))}{w}}=e^{\frac{\ell_{2}(g(w))}{w}}.$
@@
(10.1) $\ell_{1}(g(w))-\ell_{2}(g(w))=2n\pi iw.$
But $\ell_{1}(g(w))-\ell_{2}(g(w))$ is an integer multiple of $2\pi i$ for any $w\in D^{+}\cup D^{-}$. So, (10.1) holds only for $n=0$. Hence we get $\ell_{1}(g(w))=\ell_{2}(g(w))~{}\text{for all}~{}w\in D^{+}\cup D^{-}$. This implies that $\ell_{2}$ is a branch of logarithm on $\mathbb{C}^{*}$. This contradicts the fact that there is no branch of logarithm on $\mathbb{C}^{*}$ and completes the proof. ∎
We note that the above retract is however, straightenable and thereby conjugate to a graph – we say that a complex submanifold $M$ of $\mathbb{C}^{N}$ of dimension $k$ is straightenable or rectifiable if we can apply an automorphism of $\mathbb{C}^{N}$ which maps $M$ onto $\mathbb{C}^{k}$. Getting examples of retracts which are not conjugate to graphs is more non-trivial but it is not difficult to derive it from the work of Forstneric – Globevnik – Rosay [Forstn]. To briefly recall their main result, let us recall as they do in the introduction of [Forstn] that Abhyankar and Moh in [Abynkr] and M.Suzuki in [Suzuki] proved that if $P$ is a polynomial embedding of $\mathbb{C}$ into $\mathbb{C}^{2}$, then there exists a polynomial automorphism $\psi$ of $\mathbb{C}^{2}$ such that $\psi\circ P(\mathbb{C})=\mathbb{C}\times\{0\}$. This result is not true for holomorphic embeddings of $\mathbb{C}$. By proposition $1$ in [Forstn], there exists a proper holomorphic embedding $H:\mathbb{C}\to\mathbb{C}^{N}$ such that for no automorphism $\psi$ of $\mathbb{C}^{N}$, $\psi\circ H(\mathbb{C})=\mathbb{C}\times\{0\}\subset\mathbb{C}^{N}$. We want to prove that ${\rm image}(H)$ is a retract of $\mathbb{C}^{N}$. Towards this the manner in which this proper holomorphic embedding $H$ is defined and attained in the proof of proposition $2$ of their paper [Forstn] is now very briefly recalled. Start with the embedding $H_{0}:\mathbb{C}\to\mathbb{C}^{N}$ defined by $H_{0}(\zeta)=(\zeta,0)$. For $j\geq 1$, $H_{j}:=\psi_{j}\circ H_{j-1}$, where $\psi_{j}$ is some automorphism of $\mathbb{C}^{N}$. These $(H_{j})$’s consitute a sequence of proper embeddings of $\mathbb{C}$ satisfying some properties. Taking their uniform limit on compacts $H:=limH_{j}$ yields the desired embedding.
Consider now, the holomorphic retraction $\rho_{0}:\mathbb{C}^{N}\to\mathbb{C}^{N}$ defined by $\rho_{0}(z,w)=(z,0)$, where $z\in\mathbb{C}$ and $w\in\mathbb{C}^{N-1}$. Note that ${\rm image}(H_{0})={\rm image}(\rho_{0})$. For $j\geq 1$, let $\rho_{j}:\mathbb{C}^{N}\to\mathbb{C}^{N}$ defined by $\rho_{j}:=\psi_{j}\circ H_{j}\circ\psi_{j}^{-1}$, where $\psi_{j}$ is the automorphism mentioned in the above construction. Observe that each $\rho_{j}$ is a holomorphic retraction with ${\rm image}(\rho_{j})={\rm image}(H_{j})$. Note that the $\rho_{j}$’s converge uniformly on compacts. If we denote the limit map by $\rho$ then, $\rho$ is also a retraction – this follows by passing to the limit in the defining equation of a retraction wherein taking a double limit is justified owing to the uniform convergence. Finally, we note that the limit of this retraction map $\rho$ satisfies ${\rm image}(H)={\rm image}(\rho)$. Hence we have verfied the existence of a one-dimensional holomorphic retract $Z$ of $\mathbb{C}^{N}$ which is not straightenable, thereby not conjugate to a graph as well.
We conclude by expositing on results about polynomial retracts of $\mathbb{C}^{2}$, the simplest of domains admitting non-trivial retracts. As mentioned in the foregoing, holomorphic retracts in $\mathbb{C}^{2}$ are not in general rectifiable. Also, it seems difficult to get a complete characterization of holomorphic retracts of $\mathbb{C}^{2}$. However it is possible to deduce a complete characterization of polynomial retracts on $\mathbb{C}^{2}$ from known results about the polynomial ring in two complex variables and this yileds in-particular, that all polynomial retracts are rectifiable. Even this deduction is not entirely trivial and needs some work which was done by Shpilrain and Yu in [Shpilrain--Yu], which we exposit upon. In order to lay this work out properly, we begin with some definitions.
@@
Let $K[x,y]$ be the polynomial algebra in two variables over a field $K$ of characteristic 00. An idempotent $K$-algebra homomorphism $\varphi$ from $K[x,y]$ into itself is called a retraction (or a projection) map on $K[x,y]$. The image of such a homomorphism is called a retract of (the polynomial ring/algebra) $K[x,y]$. Thus, a subalgebra $R$ of $K[x,y]$ is called a retract if there is an idempotent homomorphism $\varphi:K[x,y]\rightarrow$ $K[x,y]$ such that $\varphi(K[x,y])=R$.
To provide just one concrete example to illustrate the above definitions, take $p=x+x^{2}y$; then $K[p]$ is a retract of $K[x,y]$ (the retraction map here, is given by $x\mapsto p$ , $y\mapsto 0$). Of-course as always, before proceeding to such examples, one can cite the trivial ones, namely $K[x,y]$ or $(0)$; a retract of $K[x,y]$ which is neither of these two extremes, is referred to as a proper retract in this context. In [Cost], Costa proved that every proper retract of $K[x,y]$ has the form $K[p]$ for some polynomial $p\in K[x,y]$. We shall be using this result in the sequel.
It is helpful in all of what follows, to first record the elementary fact as to how polynomial retraction maps on $\mathbb{C}^{2}$ transform to (or induce) a retraction map on the polynomial ring $\mathbb{C}[z,w]$, as discussed next.
@@
A holomorphic map $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ given by $\rho(z,w)=(p(z,w),q(z,w))$ where $p,q\in\mathbb{C}[z,w]$ is a polynomial retraction if and only if there exists a retraction map $\hat{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ which take the standard generators $z,w$ to the polynomials $p(z,w)$ and $q(z,w)$.
@@
Assume that $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ is a (holomorphic) retraction map on $\mathbb{C}^{2}$ such that its components namely $p(z,w)$ , $q(z,w)$ are polynomials. Define $\hat{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ by $z\mapsto p(z,w),~{}w\mapsto q(z,w)$. Since $\rho\circ\rho=\rho$, we have $p(p(z,w),q(z,w))=p(z,w)$ and $q(p(z,w),q(z,w))=q(z,w)$. This implies that $\hat{\rho}\circ\hat{\rho}=\hat{\rho}$ i.e., $\hat{\rho}$ is a retraction map.
Conversely, assume $\hat{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ is a retraction map. Then there exists $p(z,w),\\ q(z,w)\in\mathbb{C}[z,w]$ such that $\hat{\rho}(z)=p(z,w)$ and $\hat{\rho}(w)=q(z,w)$. Let the subalgebra generated by $p(z,w)$ and $q(z,w)$, which is a retract of the polynomial algebra $\mathbb{C}[z,w]$ be denoted by $R$. Define $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ by $\rho(z,w)=(p(z,w),q(z,w))$. Since $p,q\in R$, $\hat{\rho}(p)=p$ and $\hat{\rho}(q)=q$. This implies that $p(p(z,w),q(z,w))=p(z,w)$ and $q(p(z,w),q(z,w))=q(z,w)$. But this just means that $\rho\circ\rho=\rho$ i.e., $\rho$ is a polynomial retraction mapping on $\mathbb{C}^{2}$.
@@
This is essentially an exposition of the proof of theorem $1.1$ of [Shpilrain--Yu] with greater details. Let $Z$ be a non-trivial polynomial retract of $\mathbb{C}^{2}$. There exists a retraction map $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ of the form $\rho(z,w)=(p(z,w),q(z,w))$, where $p,q\in\mathbb{C}[z,w]$ such that $\rho(\mathbb{C}^{2})=Z$. By lemma 10.4, $\rho$ induces a retraction map $\hat{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ such that $\hat{\rho}(z)=p(z,w)$ and $\hat{\rho}(w)=q(z,w)$, the components of the retraction mapping $\rho$ on $\mathbb{C}^{2}$. Note that ${\rm image}(\hat{\rho})=\mathbb{C}[p,q]$, the subalgebra generated by $p,q$ which we shall denote by $R$.
By theorem 3.5 in [Cost], $\mathbb{C}[p,q]=\mathbb{C}[r]$, for some $r=r(z,w)\in R\subset\mathbb{C}[z,w]$. Hence we may write $p=q_{1}(r),~{}q=q_{2}(r)$, for some $q_{1},q_{2}\in\mathbb{C}[z]$ i.e., $q_{1},q_{2}$ are polynomials of a single variable. Let us write out this factorization for our retraction mapping $\rho$ on $\mathbb{C}^{2}$, a bit elaborately for convenience in the computations that follow, as
$\rho(z,w)=\left(q_{1}(r(z,w)),q_{2}(r(z,w))\right).$
Suppose $q_{1}$ and $q_{2}$ have degree $n\geq 1$ and $m\geq 1$ respectively (the case where atleast one of them has degree zero i.e., is constant, is simpler and dealt with during the course of concluding the proof). It follows by the Abhyankar – Moh theorem (main theorem of [Abynkr]) that: either $n$ divides $m$ or $m$ divides $n$. Suppose without loss of generality that $\deg(q_{1})=k\cdot\deg(q_{2})$ for some integer $k\geq 1$.
Consider then the automorphism $\psi:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by $\psi(z,w)=(z-cw^{k},w)$, where $c$ is chosen so that leading term of the polynomials $q_{1}$ and $cq_{2}^{k}$ in $q_{1}-cq_{2}^{k}$ cancel out. The point of considering this polynomial automorphism $\psi$ is due to its effect in rendering a strict reduction of degree of the polynomial $q_{1}$ when we pass from our retraction map $\rho$ to its conjugate by $\psi$, explained as follows:
@@
Observe that the degree of the single variable polynomial $\tilde{q}_{1}$ is strictly less than that of $q_{1}$; we shall call the polynomials $\tilde{q}_{1},\tilde{q}_{2}$ now simply by $q_{1},q_{2}$ and repeat the above process of passing from one retraction to another through conjugation by the (new and appropriate) automorphism $\psi$. So, continuing this process , we eventually arrive within finitely many steps, at a pair of polynomials such that the following happens. One of them is a constant polynomial $c$, and the other polynomial, denoted by $h$, is such that $\mathbb{C}[h]=\mathbb{C}[r]$ (i.e. $h=c_{1}r+c_{2}$). The composition of the various (finitely many) automorphisms involved in the aforementioned process, gives a polynomial automorphism $\phi$ of $\mathbb{C}^{2}$ such that
$\left(\phi\circ\rho\circ\phi^{-1}\right)(z,w)\;=\;(c_{1}r(z,w)+c_{2},c).$
Let $\tilde{\rho}:=\phi\circ\rho\circ\phi^{-1}$, which of-course is a retraction map on $\mathbb{C}^{2}$. After conjugating it by an affine automorphism namely $\tilde{\psi}(z,w)=(z,w-c)$, we can ensure that the retraction map – which we shall call $\tilde{\rho}$ again – has a particularly simple form wherein the second component is just zero i.e., $\tilde{\rho}(z,w)$ is of the form $(s(z,w),0)$ for some $s(z,w)\in\mathbb{C}[z,w]$. Since $\tilde{\psi}\circ\tilde{\rho}\circ\tilde{\psi}^{-1}$ is a retraction map, it follows that $s(s(z,w),0)=s(z,w)$. Writing $s(z,0)=a_{0}+a_{1}z+a_{2}z^{2}+\cdots+a_{k}z^{k}$, we then have:
$s(s(z,w),0)=a_{0}+a_{1}s(z,w)+\cdots+a_{k}s(z,w)^{k}=s(z,w).$
Comparing the highest degree monomials on both sides (of the second equality above), we first deduce that: $a_{2}=\cdots=a_{k}=0$ and $a_{1}=0$; it is then immediate that $a_{0}$ must also be zero. Thus, $s(z,0)=z$ meaning that among the monomials not divisible by $w$, there are no monomials other than $z$. Hence, we conclude that $s$ must be of the form: $s(z,w)=z+w\cdot l(z,w)$, for some $l(z,w)\in\mathbb{C}[z,w]$. Therefore in all, it follows from the foregoing arguments that there exists a polynomial automorphism $\Phi$ on $\mathbb{C}^{2}$ such that
$\tilde{\rho}(z,w)\;=\;\left(\Phi\circ\rho\circ\Phi^{-1}\right)(z,w)\;=\;\left(s(z,w),0\right).$
Thus, after applying the polynomial change of coordinates given by $\Phi$, the retract $Z$ is transformed into
$\Phi(Z)=\{(\zeta,0):\zeta\in\mathbb{C}\},$
finishing the proof of the theorem that all non-trivial polynomial retracts of $\mathbb{C}^{2}$ can be rectified. ∎